CORPORATE ANALYSIS CA2.32 The Procter & Gamble Company. The 2019 annual report of The Procter & Gamble Company (P&G) is available at www.pginvestor.com. After reviewing P&G’s annual report, respond to the following questions:

a. P&G’s net income decreased from 2017 through 2019, while sales increased during the sale period. What would account for this?

b. P&G’s total assets decreased by approximately $3.2 billion from 2018 to 2019. Which assets principally accounted for this decline? (To answer this question, calculate each asset as a percentage of total assets and then compare between years.) Is this decline a problem? If so, why? If not, why not?

c. Consider the change in P&G’s total debt and total shareholders’ equity from 2018 to 2019. Did the $3.2 billion decrease in its assets have an impact on debt? What was the impact on shareholders’ equity? Calculate P&G’s long-term debt-to-equity ratio and total debt to total assets ratio for both years as part of your analysis. What can you say about P&G’s apparent financing strategy?

d. Consider P&G’s statement of cash flow. What were the five major cash outflows for the company in 2019? How did P&G finance these cash outflows?

e. Compare P&G’s dividend payments to its net income in 2018 and 2019. (Note: The ratio of dividends paid divided by net income is called the dividend payout ratio.) What percentage of net income does P&G pay to its shareholders? How does this payout percentage compare to P&G’s competitors (such as Johnson & Johnson)? Do you agree with this dividend payment policy? What message does this policy convey to investors about the company’s future growth potential
Answer #1

(figures in billions USD)

2019 2018 2017
Sales $67.7 $ 66.8 $65.10
net income $3.9 $ 9.8 $ 10

a) if you see only numbers, we could think sales increase 1.34 % while net income decrease 60% ,. but for 2018 net income represents 14.70 % while for 2019 is 5.70 %

it would be a lot of things, like increase the depreciation for fixed assets, more costs or expenses, we must analyze every account on this part., or an increase in interest paid.

b)

Amounts in millions; As of June 30
Assets 2019 % 2018
CURRENT ASSETS
Cash and cash equivalents $                       4,239 65.01% $           2,569
Available-for-sale investment securities $                       6,048 -34.83% $           9,281
Accounts receivable $                       4,951 5.66% $           4,686
INVENTORIES
Materials and supplies $                       1,289 -3.45% $           1,335
Work in process $                          615 4.59% $               588
Finished goods $                       3,116 10.69% $           2,815
Total inventories $                       5,017 $           4,738
Prepaid expenses and other current assets $                       2,218 8.41% $           2,046
TOTAL CURRENT ASSETS $                    22,473 $         23,320
PROPERTY, PLANT AND EQUIPMENT, NET $                     21,271 3.26% $         20,600
GOODWILL $                     40,273 -10.85% $         45,175
TRADEMARKS AND OTHER INTANGIBLE ASSETS, NET $                     24,215 1.31% $         23,902
OTHER NONCURRENT ASSETS $                       6,863 29.17% $           5,313
TOTAL ASSETS $                  115,095 $       118,310

the 2 account that there is an important decrease is

Available-for-sale investment securities -34.83 % that its their function, to finance the company

an the other one is Goodwill -10.35%  

This value is a compensation paid in the stocks so, its nature is to finance the company too.

there is no problem with the asssets because the working capital is supporting all the operations, it just a different stucture in the balance sheet

c) Company debt policy is to finance the company with shareholders instead of loans with more financial cost , the shareholder equity decrease because the interest payments

Long Term debt to Total Assets Ratio = Long Term Debt / Total Assets = 20,395/115,095 =.1770 20,863/118,310=.1763 , this ratio is consistent and is conservative

Long-term debt ÷ (Common stock + Preferred stock) = Long-term debt to equity ratio = 20,395/7256+2563 =2.71

20863/7057+265=2.84 this means that we are paying our long term debts from one year to another, the level of leverage is decreasing , and this is ok

d) financing 1) treasury stock purchases $ 5,003 2) dividens to share holders $ 7,498, chage to short term debt 2,215

Investing 1) net acquisitions 3,945 and 2) capital expenditures 3,347

the company finance with the sells of investment securities, and good will , and of course with the increasing sales

e)

Dividends per Common Share $2.90 / $2.79 / $2.70 / $2.66 / $2.59 the company has 129 years giving dividends, and 63 years increasing dividends so people and markets trust in the company

if you see the dividend in Johnson & Johnson they are paying .90 per share son its less than P&G